cookiemonster:

Solve: 15x^3 + 10x^2 divided by / 5x

cookiemonster:

@Elsa213 @ThisGirlPretty @Tr3 @Shadow @Zepdrix

ThisGirlPretty:

@Vocaloid

cookiemonster:

and its not 17x^3 because my teacher said that was wrong..

Nnesha:

\[\frac{ 15x^3 +10x^2}{ 5x }=\frac{ 15x^3 }{ 5x }+\frac{ 10x^2 }{ 5x }\] where x^3 can be written as \[\rm x^3= x \cdot x\cdot x\] rewrite the first term \[\frac{ 15 }{ 5}\frac{ x \cdot x\cdot x }{ x}\] \[\frac{ \color{Red}{3} \cancel{15}}{ \cancel{5}}\frac{ \cancel{x} \cdot x\cdot x }{ \cancel{x}}= 3\cdot x \cdot x = 3x^2\]

Nnesha:

simplify the 2nd term just like that^^

cookiemonster:

ok so the final answer is 3x^2?@Nnesha

Nnesha:

not the *final answer*

cookiemonster:

so how do we get the final answer

Nnesha:

that's just the first term after simplifying

cookiemonster:

ok so whats next

Nnesha:

we get the final answer after simplifying the *2nd term* so try to do that

cookiemonster:

ok but what do i have to simplify?

Nnesha:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Nnesha \[\frac{ 15x^3 +10x^2}{ 5x }=\frac{ 15x^3 }{ 5x }+\frac{ 10x^2 }{ 5x }\] where x^3 can be written as \[\rm x^3= x \cdot x\cdot x\] rewrite the first term \[\frac{ 15 }{ 5}\frac{ x \cdot x\cdot x }{ x}\] \[\frac{ \color{Red}{3} \cancel{15}}{ \cancel{5}}\frac{ \cancel{x} \cdot x\cdot x }{ \cancel{x}}= 3\cdot x \cdot x = 3x^2\] \(\color{#0cbb34}{\text{End of Quote}}\) separate the number and variable first ^like that

cookiemonster:

so now i have to simplify using the term 2?

Nnesha:

\[\frac{ 10x^2 }{ 5x}\] simplify this

cookiemonster:

ok..

cookiemonster:

2x^3

cookiemonster:

now what??

Nnesha:

how did you get x^3 ?

Nnesha:

if x^3 is same as x *x*x how would you write x^2 term ?

cookiemonster:

bc i added 2+1

cookiemonster:

so 0?

cookiemonster:

or 2x^2

Nnesha:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @cookiemonster bc i added 2+1 \(\color{#0cbb34}{\text{End of Quote}}\) 2+1 ???

cookiemonster:

3

Nnesha:

no... i mean why are you adding 2 and 1 ?

cookiemonster:

to simplify it

Nnesha:

to simplify what ?

cookiemonster:

oh nooo just kidding i did it wrong sorry i got 2x now

Nnesha:

did you get that or someone gave it to you

cookiemonster:

no i did it.. so now we put it together.. and the answer would be x(3x +2)?

cookiemonster:

@Nnesha

Nnesha:

that's one way to write the answer. so yes

cookiemonster:

ok awesome thanksss so much

cookiemonster:

how else would u write it?

cookiemonster:

@Nnesha

Nnesha:

i will not take out the x term.

cookiemonster:

yes i got it wrong my teacher said... 3x + 2x

dude:

I would write the other answer as the current one could still be simplified and usually educators ask for answers that can no longer be simplified

Nnesha:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @cookiemonster yes i got it wrong my teacher said... 3x + 2x \(\color{#0cbb34}{\text{End of Quote}}\) ur teacher said the answe is `3x+2x` ??

cookiemonster:

yes

Nnesha:

first of all i can't believe she said that. (i'm pretty sure about that) 2nd: take a screenshot of my work and email it to her/him.

cookiemonster:

lol

cookiemonster:

idk she was kinda being mean today :/

Nnesha:

nice advice to you stop lying and try to focus on your work that will save some time. thanks

cookiemonster:

are you saying this to my teacher? lol

Nnesha:

no. that was for you

cookiemonster:

why would i lie to you about what she told me...

cookiemonster:

im sad i got it wrong im not lying i really thought i was going to get it right :/

Nnesha:

waht did you put for the *final answer*?

cookiemonster:

i put x ( 3x+2).. she was like nothing goes in front... its 3x + 2x.. im like, "ummm okay?"

Nnesha:

i can tell this is a multiple choice i already simplified the first term and you got 2x for the 2nd term just put them together

cookiemonster:

it wasnt multiple choice though

Nnesha:

or in other words x ( 3x+2) distribute the parenthesis by x

cookiemonster:

okay thats fine i learned from my mistakes... i guess next time i have to be more careful.. but all in all, i do really appreciate you helping me out truly :)

dude:



cookiemonster:

ok yea that looks right thanks for correcting me